# A Benchmark Study of Deep-RL Methods for Maximum Coverage Problems over Graphs

### v1

@@ Abstract.
Recent years have witnessed a growing trend toward employing deep reinforcement learning (Deep-RL) to derive heuristics for combinatorial optimization (CO) problems on graphs. Maximum Coverage Problem (MCP) and its probabilistic variant on social networks, Influence Maximization (IM), have been particularly prominent in this line of research. In this paper, we present a comprehensive benchmark study that thoroughly investigates the effectiveness and efficiency of five recent Deep-RL methods for MCP and IM. These methods were published in top data science venues, namely S2V-DQN [1], Geometric-QN [2], GCOMB [3], RL4IM [4], and LeNSE [5]. Our findings reveal that, across various scenarios, the Lazy Greedy algorithm consistently outperforms all Deep-RL methods for MCP. In the case of IM, theoretically sound algorithms like IMM and OPIM demonstrate superior performance compared to Deep-RL methods in most scenarios. Notably, we observe an abnormal phenomenon in IM problem where Deep-RL methods slightly outperform IMM and OPIM when the influence spread nearly does not increase as the budget increases. Furthermore, our experimental results highlight common issues when applying Deep-RL methods to MCP and IM in practical settings. Finally, we discuss potential avenues for improving Deep-RL methods. Our benchmark study sheds light on potential challenges in current deep reinforcement learning research for solving combinatorial optimization problems.
@@
PVLDB, 14(1): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
@@ 1. Introduction
Combinatorial optimization problems play a pivotal role in addressing complex challenges across diverse domains. The Maximum Coverage Problem (MCP) and its probabilistic variant, Influence Maximization (IM), stand out as representative challenges in this realm. The primary goal of MCP is to identify a set $S$ of $k$ nodes from a given input graph $G$ to maximize node coverage. On the other hand, IM focuses on maximizing influence spread in social networks by selecting a strategic set of nodes. The significance of MCP and IM transcends theoretical domains, finding practical applications in scheduling [6, 7], facility location [8, 9], recommendation systems [10, 11], viral marketing [12], and sensor placement [13]. Therefore, devising effective and efficient algorithmic techniques for solving MCP and IM has drawn great attention from the data management and data mining research communities.
@@
Despite the substantial theoretical advancements in MCP and IM, recently, there has been a notable surge in the application of Deep-RL to these combinatorial optimization problems [1, 4, 2, 3, 5]. This contemporary approach seeks to harness the power of data-driven learning, aiming to derive heuristics that can surpass the performance of theoretically grounded algorithms in practice. In these studies, Graph Neural Networks (GNNs) initially learn node embeddings, followed by the integration of reinforcement learning components like Q-learning to approximate the objective function. Empirical studies of these works reveal instances where they outperform traditional algorithms for both MCP and IM.
@@
  1. (1)
    We summarize the general pipeline of recent Deep-RL methods [1, 4, 2, 3, 5] for MCP and IM in § 3. Additionally, in § 3.4 and 3.5, we articulate concerns regarding the exclusion of training time and the absence of a strong baseline, Lazy Greedy, for MCP in the Deep-RL studies, which motivate our study.
  2. (2)
    § 4 details an extensive benchmark study aimed at examining recent Deep-RL methods for MCP and IM. We successfully reproduce some of the experimental results reported in prior studies [1, 4, 2, 3, 5] and expand more. Our findings, illustrated in Fig. 1, reveal that, despite reported successes, traditional and well-established MCP and IM algorithms often outperform these Deep-RL methods in terms of both effectiveness and efficiency.
  3. (3)
    § 5 delves deeper into the practical challenges faced by Deep-RL methods. We highlight common issues, such as the difficulty in determining the suitability of a testing graph for a trained Deep-RL model and the observed performance fluctuations concerning training time and data size.
  4. (4)
    Our experimental findings prompt a discussion on the challenges that must be addressed for Deep-RL methods to provide effective and efficient solutions for MCP and IM, offering insights for future research in this domain (§ 8).
@@
For conciseness, certain evaluations, including those for GCOMB’s noise predictor, strategies to enhance LeSNE’s efficiency, and outcomes from more datasets are omitted from the main text. Readers are encouraged to refer to the Appendix.
@@ 2. Preliminaries
In this section, we formally revisits the Maximum Coverage Problem (MCP) and its probabilistic variant in social networks, Influence Maximization (IM).
@@ 2.1. Problem Statements
Independent Cascade (IC) Model. Given a weighted and directed graph $G=\langle V,E,W\rangle$, we assume that the edge weight $p_{uv}$ on an edge $(u,v)$ represents its influence probability $p_{uv}$, i.e., $0\leq p_{uv}=w_{uv}\leq 1$. An influence diffusion starts from a seed set $S\subseteq V$. $S_{i}$ denotes the set of nodes that are active in time-step $i$, $i\in\mathbb{N}$, $S_{0}=S$. Each newly-activated vertex $u$ in the previous step $i-1$ has a single chance at the current step $i$ to influence its inactive out-neighbor $v$ independently with a probability $p_{uv}$. The diffusion process continues until there are no further activations, i.e., $S_{t}=S_{t-1}$. The influence spread of $S$, $I(S)$, is the expected number of active nodes at the end of the diffusion initiated from $S$.
As IC model is the most popular influence model, we employ IC model to formulate the IM problem.
@@ 2.3. Edge Weight Models in IM
Learned (LND) Model [3]. When we have historical data about user interactions, we may learn edge weights. One representative method to learn the edge weights is the Credit Distribution Model [26].
@@ 3. Algorithms Revisited
In this section, we delve into the foundational concepts of Deep-RL and reflect upon the Deep-RL techniques and other traditional algorithms used for the MCP/IM benchmarks in our experiments.
@@ 3.2. Deep-RL Methods for MCP/IM
GCOMB [3]. Rather than unsupervised learning, GCOMB trains a GCN network [33] in a supervised manner, where the label of each node is generated by calculating its marginal cover (influence spread) based on a probabilistic greedy method. A DQN then finds a solution set that might maximize the cover (influence spread). Node pruning techniques are also adopted to remove noisy nodes to reduce the search space, which makes GCOMB scalable to large-scale graphs.
@@ 3.3. Traditional Algorithms for MCP/IM
Greedy for MCP. Greedy algorithm [12] sequentially selects a node that covers the most remaining uncovered elements. Leveraging the submodularity of coverage functions, Lazy Greedy (also known as CELF [13]) distinguishes itself by strategically minimizing computational overhead. After initial marginal gain computations for all nodes, it selectively updates and reevaluates only the top contenders in subsequent iterations. This approach retains the $(1-\frac{1}{e})$ approximation guarantee while delivering a significant speedup over the traditional greedy algorithm. (More details please refer to Appendix.)
@@
OPIM  [18]. Contrary to the offline processing algorithm IMM, OPIM stands out as an interactive online algorithm tailored for influence maximization. Users can seamlessly request a solution along with its approximation guarantee at any stage, and have the flexibility to resume the algorithm for improved outcomes. OPIM strategically chooses nodes to maximize the anticipated influence spread, all while adhering to a set budget. Leveraging a submodular function, it ensures streamlined optimization and incorporates upper bounds to boost its efficiency.
@@ 4.1. Training Time for Deep-RL Methods
LeNSE-IM [^2] The results for LeNSE are reported after efficiency optimization (Please refer to Appendix.) 96.6 2986 13606 18056 16419
@@
Table 2. Training time of each method and number of queries answered by traditional methods (Lazy Greedy for MCP and IMM for IM) on four datasets in training time
@@ 4.2. Performance on MCP
In this section, we conduct a benchmark of Normal Greedy, Lazy Greedy, S2V-DQN, GCOMB, and LeNSE in the context of the MCP. The selective results are depicted in Fig. 4, with additional results for the remaining 12 datasets provided in the Appendix.
@@
Efficiency. Fig. 4 illustrates the computational efficiency of GCOMB, which operates 1 to 2 orders of magnitude faster than S2V-DQN and over 2 orders faster than LeNSE in most scenarios. Particularly with a small budget, it surpasses the runtime of Normal Greedy by up to two orders, consistent with the findings in [3]. The runtime of GCOMB exhibits fluctuations rather than a steady increase, primarily attributed to the varying number of good nodes predicted by its node pruner  [^3] For a more comprehensive experimental analysis, refer to the detailed results in the Appendix. In contrast, LeNSE takes over 10$\times$ longer than Normal Greedy.
@@ 4.3. Performance on IM
Figure 7. (a) Comparison between RL4IM, CHANGE, and IMM over synthetic graphs under the given weight model. The coverage is an average of the results of 10 repeated experiments. (b) Comparison between Geometric-QN(G-QN) and IMM over small-scale datasets under the given weight model. The coverage is an average of the results of 20 repeated experiments.
@@
With the help of node pruning techniques, GCOMB can sometimes achieve speeds that are orders of magnitude faster than IMM. However, the runtime of GCOMB is often non-monotonic concerning the budget, indicating that the node pruner cannot guarantee a smaller search space for a small budget compared to a large budget (e.g., TV-BrightKite, CONST-Higgs, CONST-Live Journal). The instability of the node pruner leads to GCOMB being orders of magnitude slower than IMM in certain cases.(eg. WC-Youtube, TV-Live-Journal, and more cases displayed in Appendix).
@@ 5.1. Study of Graph Distribution
Table 6. Ratio of Execution Time: Advanced Graph Similarity Calculation Approaches to OPIM Query Resolution with $k=200$
@@ 6. Rating Scale for Each Solver
Table 7. Rating scale for each solver: In each metric, higher values indicate better performance. The highest values are underlined.
Based on the observed results across a wide range of datasets and settings, a rating scale is summarized in Tab. 7. Here, the metric Quality for method $f$ is defined as the average of $\{\frac{c^{(f)}_{d}}{\text{Max}({c_{d}})}|d\in\text{D}\}$, where $c$ represents the coverage value, and $d$ denotes a selected dataset within the entire set of datasets $D$. Similarly, the metric Efficiency is defined as the average of $\{\frac{\text{Max}({t_{d}})}{t^{(f)}_{d}}|d\in\text{D}\}$, where $t$ signifies the runtime. Furthermore, we delineate the metric Robustness for method $f$ as the normalized reciprocal standard deviation of its quality. Please note that the rating for Geometric-QN, due to its limited scalability, is not derived from a direct comparison with others across $D$. Instead, it is derived by comparing it specifically with IMM on the smaller datasets, as detailed in Kamarthi et al.  [2].
@@ 7. Other Related Work
Maximum Coverage and Influence Maximization Problems. MCP is a well-known NP-hard problem that has received significant attention. Various variants of MCP, such as the Budgeted Maximum Coverage Problem [46, 47, 48, 49] and the Multiple Knapsack Problem [50], have been extensively studied, finding applications in facilities location [51, 8], maximum coverage in streams [52], and information retrieval [53], etc.
@@ 8. Potential Directions for Improving Deep-RL Methods
9. Discussion on More MCP Variants
There are some other popular MCP variants, such as Weighted MCP [63], Partial Coverage Problem [64], Stochastic MCP [65], and Generalized MCP [66]. For a detailed comparison with IM, please refer to the Appendix. IM is unique due to its relevance in network-driven processes and its ability to simulate the intricate social dynamics where each node’s influence is interlinked, presenting a unique strategic complexity.
@@ Appendix C Improve LeNSE Efficiency
The original implementation of LeNSE is slow performance, requiring several days to generate a training dataset even for small graphs. In our optimization efforts, we have significantly improved the implementation, reducing the preprocessing runtime from days (¿ 72 hours) to just a few minutes (11.3 minutes), while preserving the underlying logic. In IM problem, our optimizations include utilizing reverse influence sampling techniques for calculating influence spread, and dynamically generating training subgraphs of varying quality. As for MCP, we replace the normal greedy with Lazy Greedy to boost the efficiency in preprocessing and inference phase.
Appendix D Detailed Comparison between IM and More Variants
We provide a detailed comparison between IM with other popular MCP variants in this section.

### v2

@@ Abstract.
Recent years have witnessed a growing trend toward employing deep reinforcement learning (Deep-RL) to derive heuristics for combinatorial optimization (CO) problems on graphs. Maximum Coverage Problem (MCP) and its probabilistic variant on social networks, Influence Maximization (IM), have been particularly prominent in this line of research. In this paper, we present a comprehensive benchmark study that thoroughly investigates the effectiveness and efficiency of five recent Deep-RL methods for MCP and IM. These methods were published in top data science venues, namely S2V-DQN, Geometric-QN, GCOMB, RL4IM, and LeNSE. Our findings reveal that, across various scenarios, the Lazy Greedy algorithm consistently outperforms all Deep-RL methods for MCP. In the case of IM, theoretically sound algorithms like IMM and OPIM demonstrate superior performance compared to Deep-RL methods in most scenarios. Notably, we observe an abnormal phenomenon in IM problem where Deep-RL methods slightly outperform IMM and OPIM when the influence spread nearly does not increase as the budget increases. Furthermore, our experimental results highlight common issues when applying Deep-RL methods to MCP and IM in practical settings. Finally, we discuss potential avenues for improving Deep-RL methods. Our benchmark study sheds light on potential challenges in current deep reinforcement learning research for solving combinatorial optimization problems.
@@
PVLDB, 17(11): XXX-XXX, 2024.
doi:XX.XX/XXX.XX Z. Liang and Y. Yang contribute equally. X. Ke is the corresponding author. This work was done while X. Ke worked at the National University of Singapore.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 11 ISSN 2150-8097.
@@ 1. Introduction
Combinatorial optimization problems play a pivotal role in addressing complex challenges across diverse domains. The Maximum Coverage Problem (MCP) and its probabilistic variant, Influence Maximization (IM), stand out as representative challenges in this realm. The primary goal of MCP is to identify a set $S$ of $k$ nodes from a given input graph $G$ to maximize node coverage. On the other hand, IM focuses on maximizing influence spread in social networks by selecting a strategic set of nodes. The significance of MCP and IM transcends theoretical domains, finding practical applications in scheduling [1, 2], opinion formation and election  [3], facility location [4, 5], recommendation systems [6, 7, 8], viral marketing [9, 10, 11, 12, 13], and sensor placement [14, 15]. Therefore, devising effective and efficient algorithms for solving MCP and IM has drawn great attention from the data management and data mining research communities.
@@
Despite the substantial theoretical advancements in MCP and IM, recently, there has been a notable surge in the application of Deep-RL to these combinatorial optimization problems [21, 22, 23, 24, 25, 26, 27]. This contemporary approach seeks to harness the power of data-driven learning, aiming to derive heuristics that can surpass the performance of theoretically grounded algorithms in practice. In these studies, Graph Neural Networks (GNNs) initially learn node embeddings, followed by the integration of reinforcement learning components like Q-learning to approximate the objective function. Empirical studies of these works reveal instances where they outperform traditional algorithms for both MCP and IM.
@@
(1) We summarize the general pipeline of recent Deep-RL methods [21, 22, 23, 24, 27] for MCP and IM in § 3. Additionally, in § 3.4 and 3.5, we articulate concerns regarding the exclusion of training time and the absence of a strong baseline, Lazy Greedy, for MCP in the Deep-RL studies, which motivate our study.
(2) § 4 details an extensive benchmark study aimed at examining recent Deep-RL methods for MCP and IM. We successfully reproduce some of the experimental results reported in prior studies [21, 22, 23, 24, 27] and expand more. Our findings, illustrated in Fig. 1, reveal that, despite reported successes, traditional and well-established MCP and IM algorithms often outperform these Deep-RL methods in terms of both effectiveness and efficiency.
(3) § 5 delves deeper into the practical challenges faced by Deep-RL methods. We highlight common issues, such as the difficulty in determining the suitability of a testing graph for a trained Deep-RL model and the observed performance fluctuations concerning training time and data size.
(4) Our experimental findings prompt a discussion on the challenges that must be addressed for Deep-RL methods to provide effective and efficient solutions for MCP and IM, offering insights for future research in this domain (§ 8).
For conciseness, certain details are omitted from the main text. These include the algorithms discussed, additional variants of MCP, certain evaluations such as those for GCOMB’s noise predictor, strategies to enhance LeSNE’s efficiency, and outcomes from more datasets. Readers are encouraged to refer to the appendices in our full version [33].
@@ 2. Preliminaries
In this section, we formally revisits the Maximum Coverage Problem (MCP) and its probabilistic variant Influence Maximization (IM).
@@ 2.1. Problem Statements
Independent Cascade (IC) Model. Given a weighted and directed graph $G=\langle V,E,W\rangle$, we assume that the edge weight $p_{uv}$ on an edge $(u,v)$ represents its influence probability $p_{uv}$, i.e., $0\leq p_{uv}=w_{uv}\leq 1$. An influence diffusion starts from a seed set $S\subseteq V$. $S_{i}$ denotes the set of nodes that are active in time-step $i$, $i\in\mathbb{N}$, $S_{0}=S$. Each newly-activated vertex $u$ in the previous step $i-1$ has a single chance at the current step $i$ to influence its inactive out-neighbor $v$ independently with a probability $p_{uv}$. The diffusion process continues until there are no further activations, i.e., $S_{t}=S_{t-1}$. The influence spread of $S$, $I(S)$, is the expected number of active nodes at the end of the diffusion initiated from $S$. As IC model is the most popular influence model, we employ IC model to formulate the IM problem.
@@ 2.3. Edge Weight Models in IM
Learned (LND) Model [24]. When we have historical data of user interactions, we learn edge weights. One representative method to learn the weights is the Credit Distribution Model [36].
@@ 3. Algorithms Revisited
In this section, we delve into the foundational concepts of Deep-RL and reflect upon the Deep-RL techniques and other traditional algorithms used for the MCP/IM benchmarks in our experiments, and we also raise concerns on these studies.
@@ 3.2. Deep-RL Methods for MCP/IM
GCOMB [24]. Rather than unsupervised learning, GCOMB trains a GCN network [43] in a supervised manner, where the label of each node is generated by calculating its marginal cover (influence spread) based on a probabilistic greedy method. A DQN then finds a solution set to maximize the cover (influence spread). Node pruning techniques are also adopted to remove noisy nodes to reduce the search space, which makes GCOMB scalable to large-scale graphs.
@@ 3.3. Traditional Algorithms for MCP/IM
Greedy for MCP. Greedy algorithm [9] sequentially selects a node that covers the most remaining uncovered elements. Leveraging the submodularity of coverage functions, Lazy Greedy (also known as CELF [14]) distinguishes itself by strategically minimizing computational overhead. After initial marginal gain computations for all nodes, it selectively updates and reevaluates only the top contenders in subsequent iterations. This approach retains the $(1-\frac{1}{e})$ approximation guarantee while delivering a significant speedup over the traditional greedy algorithm. (More details can be found in Appendix of our full paper [33].)
@@ 4.1. Training Time for Deep-RL Methods
Table 2. Training time of each method and number of queries answered by traditional methods (Lazy Greedy for MCP and IMM for IM) on four datasets in the same training time.
@@
LeNSE-IM [^3] The results for LeNSE are reported after efficiency optimization (Please refer to Appendix in our full paper). 96.6 2986 13606 18056 16419
@@ 4.2. Performance on MCP
In this section, we conduct a benchmark of Normal Greedy, Lazy Greedy, S2V-DQN, GCOMB, and LeNSE for MCP. The selective results are depicted in Fig. 4. Additional results for the remaining 12 datasets can be found in Appendix of our extended version [33].
@@
Efficiency. Fig. 4 illustrates the efficiency of GCOMB, which is 1 to 2 orders of magnitude faster than S2V-DQN and over 2 orders faster than LeNSE in most scenarios. Particularly with a small budget, it surpasses the runtime of Normal Greedy by up to two orders, consistent with the findings in [24]. The runtime of GCOMB exhibits fluctuations rather than a steady increase, primarily attributed to the varying number of good nodes predicted by its node pruner  [^4] For a more comprehensive experimental analysis, refer to the detailed results in the Appendix of our full paper. In contrast, LeNSE takes over 10$\times$ longer than Normal Greedy.
@@ 4.3. Performance on IM
Figure 7. (a) Comparison between RL4IM, CHANGE, and IMM over synthetic graphs. The coverage is averaged of over 10 repeated experiments. (b) Comparison between Geometric-QN(G-QN) and IMM over small-scale datasets. The coverage is an average of the results of 20 repeated experiments.
@@
With the help of node pruning techniques, GCOMB can sometimes achieve speeds that are orders of magnitude faster than IMM. However, the runtime of GCOMB is often non-monotonic concerning the budget, indicating that the node pruner cannot guarantee a smaller search space for a small budget compared to a large budget (e.g., TV-BrightKite, CONST-Higgs, CONST-Live Journal). The instability of the node pruner leads to GCOMB being orders of magnitude slower than IMM in certain cases.(eg. WC-Youtube, TV-Live-Journal, and more cases displayed in our full version).
@@ 5.1. Study of Graph Distribution
Table 6. Ratio of runtime: Advanced graph similarity calculation approaches to OPIM query resolution with $k=200$
@@ 6. Rating Scale for Each Solver
Table 7. Rating scale for each solver: Higher values indicate better performance. The highest values are underlined.
@@
Based on the observed results across a wide range of datasets and settings, a rating scale is summarized in Tab. 7. Here, the metric Quality for method $f$ is defined as the average of $\{c^{(f)}_{d}/\text{Max}({c_{d}})|d\in\text{D}\}$, where $c$ represents the coverage value, and $d$ denotes a selected dataset within the entire set of datasets $D$. Similarly, the metric Efficiency is defined as the average of $\{\frac{\text{Max}({t_{d}})}{t^{(f)}_{d}}|d\in\text{D}\}$, where $t$ signifies the runtime. Furthermore, we delineate the metric Robustness for method $f$ as the normalized reciprocal standard deviation of its quality. Please note that the rating for Geometric-QN, due to its limited scalability, is not derived from a direct comparison with others across $D$. Instead, it is derived by comparing it specifically with IMM on the smaller datasets, as detailed in Kamarthi et al.  [23].
@@ 7. Other Related Work
Maximum Coverage and Influence Maximization Problems. MCP is a well-known NP-hard problem that has received significant attention. Various variants of MCP, such as the Budgeted Maximum Coverage Problem [55, 56, 57, 58] and the Multiple Knapsack Problem [59], have been extensively studied, finding applications in facilities location [60, 4], maximum coverage in streams [61], social recommendation [62, 63] and information retrieval [64], etc.
@@ 9. Conclusions
Acknowledgements.
This work was supported in part by the Hong Kong Research Grants Council (ECS 21214720), Alibaba Group through Alibaba Innovative Research (AIR) Program, Zhejiang Province’s “Lingyan” R&D Project under Grant No. 2024C01259, the Ningbo Yongjiang Talent Introduction Programme (2022A-237-G), the Ministry of Education, Singapore, under its MOE AcRF TIER 3 Grant (MOE-MOET32022-0001). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the funding agencies.
@@ Appendix C Improve LeNSE Efficiency
The original implementation of LeNSE is slow performance, requiring several days to generate a training dataset even for small graphs. In our optimization efforts, we have significantly improved the implementation, reducing the preprocessing runtime from days (¿ 72 hours) to just a few minutes (11.3 minutes), while preserving the underlying logic. In IM problem, our optimizations include utilizing reverse influence sampling techniques for calculating influence spread, and dynamically generating training subgraphs of varying quality. As for MCP, we replace the normal greedy with Lazy Greedy to boost the efficiency during preprocessing and inference.
Appendix D Discussion on More MCP Variants
There are some other popular MCP variants, such as Weighted MCP [74], Partial Coverage Problem [75], Stochastic MCP [76], and Generalized MCP [77]. IM is unique due to its relevance in network-driven processes and its ability to simulate the intricate social dynamics where each node’s influence is interlinked, presenting a unique strategic complexity.
We provide a detailed comparison between IM with other popular MCP variants as follow: